solve the following simultaneous equation
15x-7y=14 (1/4)
5x-y=3 (3/4)


Answer by Alan3354(69443)

solve the following simultaneous equation
15x-7y=14 (1/4)
5x-y=3 (3/4)
==================
Multiply the 2nd eqn by 3, then subtract